# RestoreGrad: Signal Restoration Using Conditional Denoising Diffusion Models with Jointly Learned Prior

### v1

@@ Abstract
Denoising diffusion probabilistic models (DDPMs) can be utilized for recovering a clean signal from its degraded observation(s) by conditioning the model on the degraded signal. The degraded signals are themselves contaminated versions of the clean signals; due to this correlation, they may encompass certain useful information about the target clean data distribution. However, existing adoption of the standard Gaussian as the prior distribution in turn discards such information, resulting in sub-optimal performance. In this paper, we propose to improve conditional DDPMs for signal restoration by leveraging a more informative prior that is jointly learned with the diffusion model. The proposed framework, called RestoreGrad, seamlessly integrates DDPMs into the variational autoencoder framework and exploits the correlation between the degraded and clean signals to encode a better diffusion prior. On speech and image restoration tasks, we show that RestoreGrad demonstrates faster convergence (5-10 times fewer training steps) to achieve better quality of restored signals over existing DDPM baselines, and improved robustness to using fewer sampling steps in inference time (2-2.5 times fewer), advocating the advantages of leveraging jointly learned prior for efficiency improvements in the diffusion process.
@@ 1 Introduction
However, for the diffusion model to adequately learn the reverse process, a large number of training iterations may be required, leading to potentially slow model convergence. Such inefficiency was recently related to the discrepancy between the real data distribution and the accustomed choice of the standard Gaussian prior by Lee et al. (2022). They have thus proposed a simple yet effective approach called PriorGrad, whose main idea is to construct a better prior by using some rule-based approaches to extract useful information from the conditioner data. However, despite improving performance on some generative speech tasks, handcrafting a “better” prior requires certain knowledge about the data characteristics, and such guidance may not always exist.
@@
To explore the idea, we introduce RestoreGrad, a new paradigm for improving conditional DDPM by jointly learning the prior distribution with the diffusion model, focusing on signal restoration applications. We apply RestoreGrad to speech enhancement (SE) and image restoration (IR) tasks to demonstrate its generality for signals of different nature. For SE, we compare with PriorGrad (Lee et al., 2022) which provides guidance on handcrafting suitable priors in the speech domain. For IR, we show that RestoreGrad serves as a promising solution for improving the baseline DDPM even in a domain that lacks such a recipe for engineering the prior. As shown in Figure 2, models trained using RestoreGrad are more data and compute-efficient than the baseline DDPM and PriorGrad; they converge faster to achieve higher quality of the restored signal. Further shown in Figure 3, the learned prior is more informative as it better correlates with the desired signal than an isotropic covariance, potentially simplifying the diffusion trajectory for improved model efficiency.
@@
Figure 2: Model learning performance. (Top) In speech domain, RestoreGrad outperforms PriorGrad (Lee et al., 2022), a recently proposed improvement to baseline DDPM (CDiffuSE) by leveraging handcrafted priors. (Bottom) In image domain, RestoreGrad provides a paradigm to improve DDPM baseline (RainDropDiff) where there is no existing recipe for computing better priors.
Figure 3: Visualizing the learned prior distribution. Here, the assumed prior form is: $p_{\psi}(\boldsymbol{\epsilon}|\mathbf{y})\coloneq\mathcal{N}(\boldsymbol{\epsilon};\mathbf{0},\mbox{diag}\{\boldsymbol{\sigma}^{2}_{\text{prior}}(\mathbf{y};\psi)\})$, where $\boldsymbol{\sigma}_{\text{prior}}$ is estimated by the prior encoder $\psi$ with input $\mathbf{y}$. It appears that $\boldsymbol{\sigma}_{\text{prior}}$ follows the level variation of the speech waveform and preserves the structure of the original image. This indicates that an informative prior approximating the data distribution has been obtained for improved efficiency of the diffusion process.
@@ 2 Background on DDPMs
Forward Process: DDPMs (Ho et al., 2020; Sohl-Dickstein et al., 2015) slowly corrupt the training data using Gaussian noise in the forward process. Let $q_{\text{data}}(\mathbf{x}_{0})$ be the data density of the original data $\mathbf{x}_{0}$. The forward process is a fixed Markov Chain that sequentially corrupts the data $\mathbf{x}_{0}\sim q_{\text{data}}(\mathbf{x}_{0})$ in $T$ diffusion steps, by injecting Gaussian noise according to a variance schedule $\{\beta_{t}\}_{t=1}^{T}\in[0,1)$:
@@
Reverse Process: One can generate new data samples from $q_{\text{data}}(\mathbf{x}_{0})$ by reversing the predefined forward process utilizing the same functional form. That is, we can progressively transform a noise $\mathbf{x}_{T}\sim p(\mathbf{x}_{T})=\mathcal{N}(\mathbf{x}_{T};\mathbf{0},\mathbf{I})$ back into the data by approximating the reverse of the forward transition probability. This process is defined by the joint distribution $p_{\theta}(\mathbf{x}_{0:T})$ of a Markov Chain with learned transitions:
@@
DDPM Learning Framework: Ideally, we would train the model $\theta$ with a maximum likelihood objective such that $p_{\theta}(\mathbf{x}_{0})$ is as large as possible, which is unfortunately intractable. To circumvent such difficulty, DDPMs (Ho et al., 2020) instead maximize an ELBO of the data log-likelihood, by introducing a sequence of hidden variables $\mathbf{x}_{1:T}$ and the approximate variational distribution $q(\mathbf{x}_{1:T}|\mathbf{x}_{0})$:
$\log p_{\theta}(\mathbf{x}_{0})\geq\mathbb{E}_{q(\mathbf{x}_{1:T}|\mathbf{x}_{0})}\left[\log\frac{p_{\theta}({\mathbf{x}_{0:T}})}{q(\mathbf{x}_{1:T}|\mathbf{x}_{0})}\right].$ (3)
@@
Signal Restoration by Conditional DDPMs: Signal restoration problems are concerned with recovering the original signals from their degraded observations, which are of paramount importance in reality while remaining challenging, as noises are ubiquitous and may be strong enough to cause significant degradation of the signal quality. Recently, adoption of deep generative models (Kingma & Welling, 2014; Goodfellow et al., 2014; Ho et al., 2020) for signal restoration tasks has considerably increased due to their remarkable capabilities of generating missing components in the data, with conditional DDPMs (Croitoru et al., 2023; Cao et al., 2024) demonstrating substantial promise. More formally, let $\mathbf{y}$ denote the degraded observation of the clean signal $\mathbf{x}_{0}$. Recovering $\mathbf{x}_{0}$ given $\mathbf{y}$ by a model $\theta$ can be cast as maximizing the conditional likelihood of data $p_{\theta}(\mathbf{x}_{0}|\mathbf{y})$. The problem is in general intractable, but can be approximated by using a DDPM conditioned on $\mathbf{y}$. The main idea is, without modifying the forward diffusion process (1), to learn a conditional diffusion model $\theta$ with $\mathbf{y}$ provided as input to the reverse process (Özdenizci & Legenstein, 2023):
@@
where $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{y})\coloneq\mathcal{N}(\mathbf{x}_{t-1};\boldsymbol{\mu}_{\theta}(\mathbf{x}_{t},\mathbf{y},t),\sigma_{t}^{2}\mathbf{I})$. Typically, a fixed covariance is assumed and the noise estimator network $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},\mathbf{y},t)$ is used to predict the mean.
@@ 3 Proposed Method: Integrating DDPM and VAE for Learnable Diffusion Prior
The reconstruction and prior matching terms are typically realized by an encoder-decoder architecture with $\boldsymbol{\epsilon}$ being the bottleneck representation sampled from the latent distribution. VAEs generally benefit from learnable latent spaces for good modeling efficiency. However, their generation capabilities often lag behind DDPMs that employ an iterative, more sophisticated decoding (reconstruction) process.
In this work, our aim is to embrace the best of both worlds, i.e., remarkable generation ability (DDPM) and modeling efficiency (VAE) to achieve improved output signal quality and training/sampling efficiency simultaneously. To this end, we introduce the following lower bound:
@@
By introducing a sequence of hidden variables $\mathbf{x}_{1:T}$, under the setup of conditional diffusion models where the Markov Chain assumption is employed on the forward process $q(\mathbf{x}_{1:T}|\mathbf{x}_{0})\coloneq\prod_{t=1}^{T}q(\mathbf{x}_{t}|\mathbf{x}_{t-1})$ and the reverse process $p_{\theta}(\mathbf{x}_{0:T}|\mathbf{y})\coloneq p(\mathbf{x}_{T})\prod_{t=1}^{T}p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{y})$ parameterized by a DDPM $\theta$, and assuming that $\mathbf{x}_{T}=\boldsymbol{\epsilon}$ (i.e., the latent noise of DDPM samples from the VAE latent distribution), we have the lower bound on $\log p(\mathbf{x}_{0}|\mathbf{y},\boldsymbol{\epsilon})$ in the reconstruction term of the VAE as:
$\log p_{\theta}(\mathbf{x}_{0}|\mathbf{y},\boldsymbol{\epsilon})\geq\mathbb{E}_{q(\mathbf{x}_{1:T}|\mathbf{x}_{0})}\left[\log\frac{p_{\theta}({\mathbf{x}_{0:T}}|\mathbf{y})}{q(\mathbf{x}_{1:T}|\mathbf{x}_{0})}\right],$ (8)
which is the ELBO of the conditional DDPM.
The proof (based on Markov Chain property) is provided in Appendix A. Proposition 3.1 suggests a seamless integration of the DDPM into the VAE framework as the decoder module for improved generation capabilities.
Having incorporated the DDPM as the decoder, we now discuss the encoder part (the prior matching term in (7)) of the framework. A straightforward design could be using a network $\psi$ (Prior Net) to parameterize the prior distribution as $p_{\psi}(\boldsymbol{\epsilon}|\mathbf{y})$, while assuming the posterior to be a fixed form of distribution like the standard Gaussian. However, this may in turn discard any useful information from between $\mathbf{x}_{0}$ and $\mathbf{y}$. To take advantage of the adequate correlation of $\mathbf{x}_{0}$ and $\mathbf{y}$ as in signal restoration applications, we propose to also parameterize the posterior distribution with another network $\phi$ (Posterior Net), to incorporate richer information about the target signal distribution into the learning of the prior. Together with (8), we obtain the new lower bound of the conditional data log-likelihood:
@@
The two-encoder design is inspired by Kohl et al. (2018) for image segmentation with traditional U-Nets. Here, we adopt the idea in the context of DDPM for signal restoration. Note that the Posterior Net is exclusively used in training to help the Prior Net learn a more informative prior.
Based on (9), we introduce the modified ELBO for the training objective of RestoreGrad:
Figure 4: Proposed RestoreGrad. During training, the conditional DDPM $\theta$, Prior Net $\psi$, and Posterior Net $\phi$ are jointly optimized by (11). During inference, the DDPM $\theta$ samples the latent noise $\boldsymbol{\epsilon}$ from the jointly learned prior distribution to synthesize the clean signal. (We summarize the algorithm details in Appendix C.1.)
@@
Assume the prior and posterior distributions are both zero-mean Gaussian, parameterized as $p_{\psi}(\boldsymbol{\epsilon}|\mathbf{y})=\mathcal{N}(\boldsymbol{\epsilon};\mathbf{0},\boldsymbol{\Sigma}_{\text{prior}}(\mathbf{y};\psi))$ and $q_{\phi}(\boldsymbol{\epsilon}|\mathbf{x}_{0},\mathbf{y})=\mathcal{N}(\boldsymbol{\epsilon};\mathbf{0},\boldsymbol{\Sigma}_{\text{post}}(\mathbf{x}_{0},\mathbf{y};\phi))$, respectively, where the covariances are estimated by the Prior Net $\psi$ (taking $\mathbf{y}$ as input) and Posterior Net $\phi$ (taking both $\mathbf{x}_{0}$ and $\mathbf{y}$ as input). Let us simply use $\boldsymbol{\Sigma}_{\text{prior}}$ and $\boldsymbol{\Sigma}_{\text{post}}$ hereafter to refer to $\boldsymbol{\Sigma}_{\text{prior}}(\mathbf{y};\psi)$ and $\boldsymbol{\Sigma}_{\text{post}}(\mathbf{x}_{0},\mathbf{y};\phi)$ for concise notation. Then, with the direct sampling property in the forward path $\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$ at arbitrary timestep $t$ where $\boldsymbol{\epsilon}\sim q_{\phi}(\boldsymbol{\epsilon}|\mathbf{x}_{0},\mathbf{y})$, and assuming the reverse process has the same covariance as the true forward process posterior conditioned on $\mathbf{x}_{0}$, by utilizing the conditional DDPM $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},\mathbf{y},t)$ as the noise estimator of the true noise $\boldsymbol{\epsilon}$, we have the modified ELBO associated with (9):
$-ELBO$ $=\underbrace{\frac{\bar{\alpha}_{T}}{2}\mathbb{E}_{\mathbf{x}_{0}}\lVert\mathbf{x}_{0}\rVert^{2}_{\boldsymbol{\Sigma}^{-1}_{\text{post}}}+\frac{1}{2}\log\lvert\boldsymbol{\Sigma}_{\text{post}}\rvert}_{\text{Latent Regularization (LR) terms}}$ (10)
@@
The derivation (see Appendix A) is based on combining the conditional VAE and the results in Lee et al. (2022). Notably, we join the conditional DDPM with the posterior/prior encoders and optimize all modules at once, by connecting the DDPM prior space with the latent space estimated by the encoders as in Proposition 3.1. To this end, the sampling $\boldsymbol{\epsilon}\sim q_{\phi}(\boldsymbol{\epsilon}|\mathbf{x}_{0},\mathbf{y})$ is performed by the standard reparameterization trick as in VAEs, unlocking end-to-end training via gradient descent on the obtained loss terms:
  • Latent Regularization (LR) terms: help learn a reasonable latent space; e.g., minimizing $\log\lvert\boldsymbol{\Sigma}_{\text{post}}\rvert$ avoids $\boldsymbol{\Sigma}_{\text{post}}$ from becoming arbitrary large due to the presence of its inverse in the weighted norms.
  • Denoising Matching (DM) terms: responsible for training the DDPM to predict the prior noise.
  • Prior Matching (PM) terms: obtain a desirable latent space by aligning the prior and posterior distributions.
Training of RestoreGrad: With the the conditional DDPM $\theta$, Prior Net $\psi$, and Posterior Net $\phi$ defined in Proposition 3.2, we are ready to perform optimization on learning the model parameters of $\theta,\psi,\phi$ based on the modified ELBO. The RestoreGrad framework jointly trains the three neural network modules by minimizing (10) as depicted in Figure 4. Following existing DDPM literature, we approximate the objective by dropping the weighting constant $\gamma_{t}$ of the DM terms, leading to the simplified loss:

### v3

@@ Abstract
Denoising diffusion probabilistic models (DDPMs) can be utilized to recover a clean signal from its degraded observation(s) by conditioning the model on the degraded signal. The degraded signals are themselves contaminated versions of the clean signals; due to this correlation, they may encompass certain useful information about the target clean data distribution. However, existing adoption of the standard Gaussian as the prior distribution in turn discards such information when shaping the prior, resulting in sub-optimal performance. In this paper, we propose to improve conditional DDPMs for signal restoration by leveraging a more informative prior that is jointly learned with the diffusion model. The proposed framework, called RestoreGrad, seamlessly integrates DDPMs into the variational autoencoder (VAE) framework, taking advantage of the correlation between the degraded and clean signals to encode a better diffusion prior. On speech and image restoration tasks, we show that RestoreGrad demonstrates faster convergence (5-10 times fewer training steps) to achieve better quality of restored signals over existing DDPM baselines and improved robustness to using fewer sampling steps in inference time (2-2.5 times fewer), advocating the advantages of leveraging jointly learned prior for efficiency improvements in the diffusion process.
@@ 1 Introduction
However, for the diffusion model to adequately learn the reverse process, a large number of training iterations may be required, leading to potentially slow convergence. Such inefficiency was related to the discrepancy between the real data distribution and the accustomed choice of the standard Gaussian prior by Lee et al. (2022). They therefor proposed a simple yet effective approach called PriorGrad, aiming to construct a better prior by using rule-based approaches to extract useful information from the conditioner data. However, despite improving performance on some generative speech tasks, handcrafting a “better” prior requires certain knowledge about the data characteristics, and such guidance may not always exist for a given task or application.
@@
To explore the idea, we introduce RestoreGrad, a new paradigm for improving conditional DDPM by learning the prior distribution in tandem with the diffusion model, focusing on signal restoration applications. We apply RestoreGrad to speech enhancement (SE) and image restoration (IR) tasks to demonstrate its generality for signals of different nature. For SE, we compare with PriorGrad (Lee et al., 2022) which provides guidance on handcrafting suitable priors in the speech domain. For IR, we show that RestoreGrad serves as a promising solution for improving the baseline DDPM even in a domain that lacks such a recipe for engineering the prior. As shown in Figure 2, models trained using RestoreGrad are more data and compute-efficient than the baseline DDPMs and PriorGrad; they converge faster to achieve higher quality of the restored signal. Further shown in Figure 3, the learned prior is more informative as it better correlates with the desired signal than an isotropic covariance, potentially simplifying the diffusion trajectory for improved model efficiency.
@@
Figure 2: Model learning performance. (Top) In the speech domain, RestoreGrad outperforms PriorGrad (Lee et al., 2022), a recently proposed improvement to the baseline DDPM (CDiffuSE (Lu et al., 2022)) which leverages handcrafted priors. (Bottom) In the image domain, RestoreGrad provides a paradigm to improve the DDPM baseline (RainDropDiff (Özdenizci & Legenstein, 2023)) when there is no existing recipe for obtaining better priors.
Figure 3: Visualization of the learned prior distribution. Here, the distribution of the prior is modeled as: $p_{\psi}(\boldsymbol{\epsilon}|\mathbf{y})\coloneq\mathcal{N}(\boldsymbol{\epsilon};\mathbf{0},\mbox{diag}\{\boldsymbol{\sigma}^{2}_{\text{prior}}(\mathbf{y};\psi)\})$, where $\boldsymbol{\sigma}_{\text{prior}}$ is estimated by the prior encoder $\psi$ with input $\mathbf{y}$. It appears that $\boldsymbol{\sigma}_{\text{prior}}$ follows the level variation of the speech waveform (in SE) and preserves the structure of the original image (in IR). This indicates that an informative prior approximating the data distribution has been obtained, leading to improved efficiency of the diffusion process.
@@ 2 Background on DDPMs
2.1 Forward Process
DDPMs (Ho et al., 2020; Sohl-Dickstein et al., 2015) slowly corrupt the training data using Gaussian noise in the forward process. Let $q_{\text{data}}(\mathbf{x}_{0})$ be the data density of the original data $\mathbf{x}_{0}$. The forward process is a fixed Markov Chain that sequentially corrupts the data $\mathbf{x}_{0}\sim q_{\text{data}}(\mathbf{x}_{0})$ in $T$ diffusion steps, by injecting Gaussian noise according to a variance schedule $\{\beta_{t}\}_{t=1}^{T}\in(0,1)$:
@@ 2.1 Forward Process
2.2 Reverse Process
One can generate new data samples from $q_{\text{data}}(\mathbf{x}_{0})$ by reversing the predefined forward process utilizing the same functional form. That is, we can progressively transform a noise $\mathbf{x}_{T}\sim p(\mathbf{x}_{T})=\mathcal{N}(\mathbf{x}_{T};\mathbf{0},\mathbf{I})$ back into the data by approximating the reverse of the forward transition probability. This process is defined by the joint distribution $p_{\theta}(\mathbf{x}_{0:T})$ of a Markov Chain with learned transitions:
@@ 2.2 Reverse Process
2.3 DDPM Learning Framework
Ideally, we would train the model $\theta$ with a maximum likelihood objective such that the probability assigned by the model $p_{\theta}(\mathbf{x}_{0})$ to each training example is as large as possible, which is unfortunately intractable (Croitoru et al., 2023). To circumvent such difficulty, DDPMs (Ho et al., 2020) instead maximize an ELBO of the data log-likelihood, by introducing a sequence of hidden variables $\mathbf{x}_{1:T}$ and the approximate variational distribution $q(\mathbf{x}_{1:T}|\mathbf{x}_{0})$:
$\log p(\mathbf{x}_{0})\geq\mathbb{E}_{q(\mathbf{x}_{1:T}|\mathbf{x}_{0})}\left[\log\frac{p_{\theta}({\mathbf{x}_{0:T}})}{q(\mathbf{x}_{1:T}|\mathbf{x}_{0})}\right].$ (3)
@@ 2.3 DDPM Learning Framework
2.4 Signal Restoration by Conditional DDPMs
Signal restoration is concerned with recovering the original signals from their degraded observations, which are of paramount importance in reality but remaining challenging, as noises are ubiquitous and may be strong enough to cause significant degradation of the signal quality. Recently, adoption of deep generative models (Kingma & Welling, 2014; Goodfellow et al., 2014; Ho et al., 2020) for signal restoration tasks has considerably increased due to their remarkable capabilities of generating missing components in the data, with conditional DDPMs (Croitoru et al., 2023; Cao et al., 2024) demonstrating substantial promise. More formally, let $\mathbf{y}$ denote the degraded observation of the clean signal $\mathbf{x}_{0}$. Recovering $\mathbf{x}_{0}$ given $\mathbf{y}$ by a model $\theta$ can be cast as maximizing the conditional likelihood of data $p_{\theta}(\mathbf{x}_{0}|\mathbf{y})$. The problem is in general intractable, but can be approximated by using a DDPM conditioned on $\mathbf{y}$. The main idea is to learn a diffusion model $\theta$ with $\mathbf{y}$ provided as a conditioner in the reverse process (2):
@@ 2.4 Signal Restoration by Conditional DDPMs
where $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{y})\coloneq\mathcal{N}(\mathbf{x}_{t-1};\boldsymbol{\mu}_{\theta}(\mathbf{x}_{t},\mathbf{y},t),\sigma_{t}^{2}\mathbf{I})$ assuming a fixed covariance. In practice, a noise estimator network $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},\mathbf{y},t)$ is adopted to predict the mean, following the practice in Ho et al. (2020).
@@ 3 Proposed Method: Integrating DDPM and VAE for Learnable Diffusion Prior
The reconstruction and prior matching terms are typically realized by an encoder-decoder architecture with $\boldsymbol{\epsilon}$ being the bottleneck representation sampled from the latent distribution. VAEs generally benefit from learnable latent spaces for good modeling efficiency. However, their generative performance often lags behind DDPMs that employ an iterative, more sophisticated decoding (reconstruction) process.
In this work, our aim is to embrace the best of both worlds, i.e., remarkable generative power (DDPM) and modeling efficiency (VAE) to achieve improved output signal quality and training/sampling efficiency simultaneously.
@@
By introducing a sequence of hidden variables $\mathbf{x}_{1:T}$, under the setup of conditional diffusion models where the Markov Chain assumption is employed on the forward process $q(\mathbf{x}_{1:T}|\mathbf{x}_{0})\coloneq\prod_{t=1}^{T}q(\mathbf{x}_{t}|\mathbf{x}_{t-1})$ and the reverse process $p_{\theta}(\mathbf{x}_{0:T}|\mathbf{y})\coloneq p(\mathbf{x}_{T})\prod_{t=1}^{T}p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{y})$ parameterized by a DDPM $\theta$, and assuming that $\mathbf{x}_{T}=\boldsymbol{\epsilon}$ (i.e., the latent noise of DDPM samples from the VAE latent distribution), we have the lower bound on $\log p(\mathbf{x}_{0}|\mathbf{y},\boldsymbol{\epsilon})$ in the reconstruction term of the VAE (7) as:
$\log p(\mathbf{x}_{0}|\mathbf{y},\boldsymbol{\epsilon})\geq\mathbb{E}_{q(\mathbf{x}_{1:T}|\mathbf{x}_{0})}\left[\log\frac{p_{\theta}({\mathbf{x}_{0:T}}|\mathbf{y})}{q(\mathbf{x}_{1:T}|\mathbf{x}_{0})}\right],$ (8)
which is the ELBO of the conditional DDPM (i.e., the conditional version of (3)).
The proof (based on Markov Chain property) is provided in Appendix A. Note that the assumption $\mathbf{x}_{T}=\boldsymbol{\epsilon}$ follows the standard DDPM to sample $\mathbf{x}_{T}$ from the distribution of the prior noise $\boldsymbol{\epsilon}$, practically achieved by using a large enough $T$ and a carefully designed variance schedule $\{\beta_{t}\}_{t=1}^{T}$. In our case, we adopt the same assumption to enable sampling $\mathbf{x}_{T}$ from the latent space of VAE. We interpret Proposition 3.1 as a seamless integration of DDPM into the VAE framework to achieve improved generative (decoding) capabilities.
Having incorporated the DDPM as the decoder, we now discuss the encoding part, i.e., the prior matching term in (7). A straightforward design could be using a network $\psi$ (Prior Net) to parameterize the prior distribution as $p_{\psi}(\boldsymbol{\epsilon}|\mathbf{y})$, while assuming the posterior to be a fixed form of distribution like the standard Gaussian. However, this may in turn discard any useful information inherent between $\mathbf{x}_{0}$ and $\mathbf{y}$. To take advantage of the adequate correlation present in signal restoration settings, we propose to also parameterize the posterior distribution with another network $\phi$ (Posterior Net), to incorporate richer information about the target signal distribution into the learning of the prior. Together with (8), we introduce the new lower bound of the conditional data log-likelihood:
@@
Based on the assumption $\mathbf{x}_{T}=\boldsymbol{\epsilon}$, the conditional DDPM samples the latent noise $\mathbf{x}_{T}$ from the distribution of $\boldsymbol{\epsilon}$ which is jointly estimated by the two encoders, $\phi$ and $\psi$. The two-encoder design is inspired by Kohl et al. (2018) for image segmentation with traditional U-Nets. Here, we adopt the idea in the context of DDPM for signal restoration, which is effective as it incorporates posterior information exploiting the correlation between clean and degraded signals. Based on (9), we introduce the training objective of RestoreGrad:
Figure 4: Proposed RestoreGrad. During training, the conditional DDPM $\theta$, Prior Net $\psi$, and Posterior Net $\phi$ are jointly optimized by (11). During inference, the DDPM $\theta$ samples the latent noise $\boldsymbol{\epsilon}$ from the jointly learned prior distribution to synthesize the clean signal. (Summary of the algorithm details is presented in Appendix C.1.)
@@
Assume the prior and posterior distributions are both zero-mean Gaussian, parameterized as $p_{\psi}(\boldsymbol{\epsilon}|\mathbf{y})=\mathcal{N}(\boldsymbol{\epsilon};\mathbf{0},\boldsymbol{\Sigma}_{\text{prior}}(\mathbf{y};\psi))$ and $q_{\phi}(\boldsymbol{\epsilon}|\mathbf{x}_{0},\mathbf{y})=\mathcal{N}(\boldsymbol{\epsilon};\mathbf{0},\boldsymbol{\Sigma}_{\text{post}}(\mathbf{x}_{0},\mathbf{y};\phi))$, respectively, where the covariances are estimated by the Prior Net $\psi$ (taking $\mathbf{y}$ as input) and Posterior Net $\phi$ (taking both $\mathbf{x}_{0}$ and $\mathbf{y}$ as input). Let us simply use $\boldsymbol{\Sigma}_{\text{prior}}$ and $\boldsymbol{\Sigma}_{\text{post}}$ hereafter to refer to $\boldsymbol{\Sigma}_{\text{prior}}(\mathbf{y};\psi)$ and $\boldsymbol{\Sigma}_{\text{post}}(\mathbf{x}_{0},\mathbf{y};\phi)$ for concise notation. Then, with the direct sampling property in the forward path $\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$ at arbitrary timestep $t$ where $\boldsymbol{\epsilon}\sim q_{\phi}(\boldsymbol{\epsilon}|\mathbf{x}_{0},\mathbf{y})$, and assuming the reverse process has the same covariance as the true forward process posterior conditioned on $\mathbf{x}_{0}$, by utilizing the conditional DDPM $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},\mathbf{y},t)$ as the noise estimator of the true noise $\boldsymbol{\epsilon}$, we have the modified ELBO, $-\mathcal{L}(\theta,\phi,\psi)$, associated with (9):
$\mathcal{L}(\theta,\phi,\psi)$ $=\underbrace{\frac{\bar{\alpha}_{T}}{2}\mathbb{E}_{\mathbf{x}_{0}}\lVert\mathbf{x}_{0}\rVert^{2}_{\boldsymbol{\Sigma}^{-1}_{\text{post}}}+\frac{1}{2}\log\lvert\boldsymbol{\Sigma}_{\text{post}}\rvert}_{\text{Latent Regularization (LR) terms}}$ (10)
@@
The derivation (see Appendix B) is based on combining VAE and the results in Lee et al. (2022). Notably, we join the conditional DDPM with the posterior/prior encoders and optimize all modules at once, by connecting the DDPM prior space with the latent space estimated by the encoders. To this end, the sampling $\boldsymbol{\epsilon}\sim q_{\phi}(\boldsymbol{\epsilon}|\mathbf{x}_{0},\mathbf{y})$ is performed by the standard reparameterization trick as in VAEs, unlocking end-to-end training via gradient descent on the loss terms:
  • Latent Regularization (LR) terms: to help learn a reasonable latent space; e.g., minimizing $\log\lvert\boldsymbol{\Sigma}_{\text{post}}\rvert$ avoids $\boldsymbol{\Sigma}_{\text{post}}$ from becoming arbitrary large due to the presence of its inverse in the weighted norms.
  • Denoising Matching (DM) terms: responsible for training the DDPM to predict the true noise.
  • Prior Matching (PM) terms: to shape a desirable latent space by aligning the prior and posterior distributions. Note that we model the distributions as zero-mean, based on that signals can be properly normalized.
3.1 Training of RestoreGrad
With the the conditional DDPM $\theta$, Prior Net $\psi$, and Posterior Net $\phi$ defined in Proposition 3.2, optimization can be performed to learn the model parameters of $\theta,\psi,\phi$ based on the modified ELBO. The RestoreGrad framework jointly trains the three neural network modules by minimizing (10) as depicted in Figure 4. Following existing DDPM literature, we approximate the objective by dropping the weighting constant $\gamma_{t}$ of the DM terms, leading to the simplified loss: